OpenStudy (anonymous):

find the 1st and 2nd and 3rd derivatives

OpenStudy (anonymous):



OpenStudy (anonymous):

i need help ...

OpenStudy (anonymous):

What differentiation laws are you familiar with that you could apply to this situation?

OpenStudy (anonymous):

is it derivative of product?

OpenStudy (anonymous):

that would be a very good place to start, yes it is

OpenStudy (anonymous):

nc .. but i dont have idea how to solve it..

OpenStudy (anonymous):

Right okay, \[\frac{ d }{ dx }(u \times v) = u*\frac{ dv }{ dx } + v*\frac{ du }{ dx }\] So in this question: \[u=\frac{ 1 }{ x } v=(x ^{2}-3x)^{12}\]